OpenStudy (rebeccapink6398):

What is the perimeter of this shape? A. 94 ft B. 96 ft C. 132 ft D. 156 ft {ps please give me time to draw}

OpenStudy (rebeccapink6398):



OpenStudy (anonymous):

132

OpenStudy (rebeccapink6398):

how? could you explain ??

OpenStudy (lulubj):

A, add 10+10+24+14+28+8=94. Thats how i used to do it and i got it right but idk

OpenStudy (anonymous):

28 + 24 +24 +10 + 10 +14 + 14 +8 = 132 Just add up every side edge

OpenStudy (anonymous):

well its a square...the top and bottom both equal 28, the sides equal 24. but you add 24(left side)+24(right side)+10(top left)+10(top right)+8(between the two 10's)+28(bottom)+14(right side of the left 10)+14(left side of the right 10)= 132

OpenStudy (anonymous):

24+28=52+10=62+10=72+8=90+4 in 14=94 idk